OpenStudy (anonymous):

a wheel, including the tyre, has a radius of 26cm. how many complete revolutions will the wheel make in travelling 1km? I have no clue how to do this any help please :)

OpenStudy (anonymous):

here the key is to find the circumference , can you do that using the given information

OpenStudy (anonymous):

to work out the circumference you got to do the radius times pi but i don't know how to do it with this :/

OpenStudy (anonymous):

the circumference of a circle is givne by : 2*pi*r

OpenStudy (anonymous):

in our case you would have to solve: 2*3.14*26

OpenStudy (anonymous):

whats 2*pi*r? and i am now lost lol

OpenStudy (anonymous):

that is the formula for the circumference of a circle, or in this case the wheel (since a wheel is shaped like a circle)

OpenStudy (anonymous):

yeah i get that bit

OpenStudy (anonymous):

basically we are gonna use this formula because we need to find out how much distance the wheel travels in one revolution. By finding the circufernece we area basically cutting the wheel down the middle and flattening out the wheel so that we can see just how much distance the wheel covers

OpenStudy (anonymous):

ok i've done that now, so do i have to do 26*pi=..../2 ?

OpenStudy (anonymous):

the formula is: 26*2*3.14. That should give you: 163.28

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

now there should be a unit attached to that, in this case it would have to be 163.28 cm

OpenStudy (anonymous):

so the wheel covers 163.28 cm in one revolution

OpenStudy (anonymous):

it sounds so simple now thank you very much :)

OpenStudy (anonymous):

now all we need to find is how many cm are in 1km

OpenStudy (anonymous):

100cm = 1 meter. 1 km = 1000 meters 1 km = 1000 * (1 meter) = 1000 * 100 cm = 100000 cm.

OpenStudy (anonymous):

Thus we see that 1km is actually equal to 100000cm

OpenStudy (anonymous):

yeah do i have to dive 100000 by 163.28

OpenStudy (anonymous):

yup, that is the last step

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you just got schooled

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

so is the answer 612.14 cm ?

OpenStudy (anonymous):

The answer should be 612.4 revolutions. If your teacher wants a whole number you can round it to 612 revolutions

OpenStudy (anonymous):

yeah thats what i got .

OpenStudy (anonymous):

:)